OpenStudy (anonymous):

Find the polynomial of lowest degree w/ only real coefficient and having the given zeros 2+i, 3

OpenStudy (anonymous):

if 2+i is one zero then 2-i is also one of the zero hence reqd polynomial is (x-3)( x-2-i)(x-2+i)

OpenStudy (anonymous):

=(x-3)* ( (x-2)^2 -i^2) =(x-3)*( (x-2)^2 +1 ) and so on ....